A flexible metal sheet, length 6.0 cm, breadth 5.0cm, average thickness 0.0833 cm has a mass of 23.0g. Calculate the volume and density of the metal sheet:

A. 2.499cm³ and 9.2gcm³
B. 30cm³ and 6.654gcm³
C. 3.645cm³ and 6.310cm³
D. 1.77cm³ and 12.994gcm³
1
Expert's answer
2020-01-06T10:15:08-0500

Volume = "6.0\u00d75.0\u00d70.0833= 2.499 cm^3"

"d=\frac{m}{V}=\frac{23.0g}{2.499 cm^3}=9.20 \frac{g}{cm^3}"

Answer: A.2.499 "cm^3" and 9.2 "\frac{g}{cm^3}"